Use the Acid-Base Table to write net equations and to determine the equilibrium constants for the Bronsted acid-base reactions that occur when aqueous solutions of the following are mixed. See the instructions for writing chemical equations given above.

Reactants Acid(1) + Base(2) = Base(1) + Acid(2) K

a) HClO4 and KF

b) HIO3 and NH3

c) HCN + KClO Click here to preview your answer.
Answer #1

HClO4(aq.) + KF(aq.) = HF(aq.) + KClO4(aq.) , Keq.= [HF(aq.)][KClO4(aq.)]/[HClO4(aq.)][KF(aq.)]

HIO3(aq.) + NH3(g) = NH4+ (aq.) + IO3-(aq.) , Keq.= [ NH4+(aq.)][IO3-(aq.)]/[HIO3(aq.)]

HCN(aq.) + KClO​​​​​​​(aq.) = KCN​​​​​​​(aq.) + HClO​​​​​​​(aq.) , Keq.=[KCN​​​​​​​(aq.)][HClO​​​​​​​(aq.)]/[HCN​​​​​​​(aq.)][KClO​​​​​​​(aq.)]